# SALSA-RL: Stability Analysis in the Latent Space of Actions for Reinforcement Learning

### v1

@@ 3.5 Transient Growth Analysis
In latent action dynamics, stability ensures the local stability of the system. However, even when the spectral radius condition $\rho(\mathbf{A}_{t})<1$ is met—ensuring eigenvalues of $\mathbf{A}_{t}$ lie within the unit circle—non-normality (i.e., $\mathbf{A}_{t}\mathbf{A}_{t}^{*}\neq\mathbf{A}_{t}^{*}\mathbf{A}_{t}$) can still induce transient growth. This transient amplification, driven by the properties of $\mathbf{A}_{t}$, can destabilize intermediate states, amplify noise, or disrupt training. Due to the local nature of our stability analysis, transient growth study becomes important for ascertaining whether a region in state-action phase space may ‘escape’ into a region of unstable action evolution.
For the latent dynamics described in Equation 4, the spectral radius $\rho(\mathbf{A}_{t})$ (the largest magnitude of the eigenvalues of $\mathbf{A}_{t}$) is first computed to confirm the system stability. Second, for cases where $\rho(\mathbf{A}_{t})<1$ but $\mathbf{A}_{t}$ is non-normal, transient growth is analyzed using the Kreiss constant $\eta(\mathbf{A}_{t})$, which measures the potential short-term amplification of the system. The Kreiss constant is defined as:
@@
where $z$ is sampled from the complex plane outside the unit circle (i.e., $|z|>1$). A high $\eta(\mathbf{A}_{t})$ suggests a higher likelihood of transient growth. In such a case, one may observe an agent ultimately performing in an unstable manner even though the starting point of a trajectory has real eigenvalues that are less than 1. Algorithm details are provided in Appendix B.
Overall, the combination of local spectral radius computation and Kreiss constant analysis provides a qualitative understanding of both the long-term stability and the transient behavior of the system, in the absence of global stability analysis. This ensures an indirect assessment of the robustness of the learned control through our proposed framework.
Floquet Analysis for Periodic Stability Floquet analysis [25] is a method specifically designed to evaluate the stability of periodic systems and their behavior over time. It quantifies the evolution of small perturbations in the latent space, capturing their growth, decay, or oscillatory behavior through Floquet exponents. This complements the spectral radius and Kreiss constant analyses by focusing on periodic stability, where transient growth alone may not capture recurring instabilities or oscillations. This analysis is applicable when the linear term in the latent dynamical system exhibits periodic trends, as seen in tasks requiring repeated oscillatory actions. An example is a pendulum failing to swing upright and instead oscillating at the bottom, where periodic stability is essential for sustained motion.
@@
In the SALSA-RL framework, $\mathbf{A}_{t}$ is directly derived from the state $\mathbf{s}_{t}$, enabling the period to be identified by analyzing state information. Floquet exponents are computed at the end of each episode to diagnose latent dynamics, revealing instability, oscillations, or unstable growth. This analysis identifies regions of instability, oscillatory behavior, or unstable growth, helping refine the control policy. Appendix B provides a step-by-step algorithm for this analysis.
@@ 4 Experiments
A range of classical RL environments are selected from OpenAI Gym [5], focusing on continuous control tasks such as Cartpole-v1 (modified for continuous action space), Pendulum-v1, LunarLanderContinuous-v2, and BipedalWalker-v3. Additionally, a modified variant of LunarLander is analyzed with a reward function emphasizing hovering over landing to enhance stability and interpretability (details in Appendix D).
4.1 Eigenvalue-based Stability Analysis
@@
Pendulum. This task involves swinging the pendulum upright with one action dimension. A latent size of $d_{h}=3$ is used for model training and eigenvalue-based stability analysis. Figure 2 shows the system’s behavior over time (red line) and the eigenvalue distribution (underlying contour).
Figure 2: Stability analysis of Pendulum control. The white line indicates the critical eigenvalue threshold of 1, while the red line represents observed state trajectory over time. The underlying contour depicts the eigenvalue distribution across potential states at each time step. Overall, the contours illustrate stability conditions over time, while the trajectory highlights how the control policy keeps states within the desired range to maintain stability, avoiding high-eigenvalue unstable regions (bounded by the white line).
@@
In the first column, the trajectory (red line) consistently avoids unstable control actions (yellow regions bounded by white lines) throughout the time domain, ensuring stability. Initially, as the pendulum swings up from the bottom ($\cos\theta,\sin\theta<0$), it approaches unstable regions very closely but successfully avoids them, reflecting effective training. Over time, the pendulum stabilizes upright ($\cos\theta\to 1$, $\sin\theta\to 0$), with the real eigenvalue converging to 0.478, indicating steady-state behavior in the stable region. In the second column, the imaginary eigenvalues remain at 0, confirming stability. While states initially approach regions with positive imaginary values, suggesting potential oscillations, they eventually shift away, guiding the system to a steady state.
@@
Appendix E further examines regions of instability, recovery, and control failures, reinforcing the framework’s stability analysis and interpretability. Overall, this analysis shows the framework policy not only stabilizes the pendulum but does so in a manner that aligns with the primary objective of keeping it upright. By maintaining eigenvalues within the stable region and avoiding oscillatory dynamics, the system ensures stability while providing interpretable insights into how the policy achieves the desired control objectives.
Cartpole. The Cartpole contains a single action moving left or right to keep the pole upright. With $d_{h}=3$, the trained model effectively controls the system, avoiding unstable regions, as shown in Figure 3. A key observation in the third row, representing the pole angle state, is that the real eigenvalue analysis identifies bounded regions above and below the central axis. These regions correspond to large angular deviations that not only lead to state instability, making stable actions or states difficult to maintain, but also aligning with the problem’s objective: exceeding these thresholds results in episode termination. This analysis interprets that the control policy not only stabilizes the Cartpole but also aligns with the primary goal of keeping the pole upright while avoiding instability-inducing angular thresholds.
Figure 3: Stability analysis of CartPole control. Three representative states are depicted in the rows. In the third row (pole angle), the trajectory and surrounding bounded regions highlight how the control policy maintains CartPole states within stable limits.
@@
LunarLander (Hovering Objective). For our lunar lander experiment, we utilize a modified environment with a hovering objective instead of an objective to land safely as typically utilized. Since the state includes a 2D space, stability is evaluated across the entire 2D domain, with each frame in Figure 4 representing a single timestep.
@@
Figure 4: Stability analysis of LunarLander hovering control. The white line marks the stability threshold at $\lambda_{1}=1$. The lander is initially deployed in various regions beyond the default upper-middle position (case 1). Each contour (left to right) shows eigenvalue distributions in 2D space, highlighting stability regions, with the trajectory depicting movement over 1,000 timesteps. Notably, case 2 contours provide early warnings of instability before the lander crashes. Animations are provided in the Supplementary.
@@
Among the four cases, the real part of the eigenvalue consistently evolves from high-value regions to lower-value regions, crossing the stability threshold (white line) and transitioning from unstable regions to stable ones. This behavior is particularly notable in cases 1, 3, and 4 during later frames or timesteps. The imaginary contour indicates oscillatory regions, but the actual lander trajectory avoids these regions in all cases. In case 2, however, the lander crashes at the final frame ($T=150$). The real eigenvalue contour in this case reveals large regions of instability, aligning with the lander’s inability to maintain its hovering status. Notably, detecting that the current action-space combination resides in a high-instability region effectively “predicts” an impending crash before it happens. This strongly underscores the value of our stability analysis, as it provides early warnings of imminent failures.
Besides the above 3 baseline experiments, Appendix F explores the BipedalWalker benchmark with additional results and discussions, reinforcing the framework and analysis.
@@
Figure 5: Pendulum control comparisons. The pendulum achieves stabilization in case 1 (first column) while displaying periodic behavior with oscillations in case 2 (second column). The Kreiss constant (third row) shows differing transient growth behaviors over time (regions of instability or normality are not evaluated).
4.2 Transient Growth Analysis
Extending the eigenvalue-based stability analysis, the Kreiss constant $\eta$ is evaluated to assess transient growth in pendulum under two scenarios: a standard setup (Case 1) and a modified environment (Case 2), where gravity increases from 10 to 15 and mass from 1.0 to 1.1. The third row of Figure 5 illustrates $\eta$ over time for both cases.
In Case 1, the pendulum successfully stabilizes in the upright position. This is reflected in the actions and states, which converge to near-zero values without oscillations (first and second rows). The Kreiss constant remains small throughout, indicating minimal transient growth and robust stability. In Case 2, the increased gravity and mass disrupt the stabilization process, causing the pendulum to exhibit periodic motion instead of reaching the upright position. This behavior is evident in the actions and states, which display sustained oscillations over time. The Kreiss constant shows periodic spikes, reflecting transient growth that aligns with the observed oscillatory response. These spikes indicate significant susceptibility to transient amplification, even though the system remains bounded over time.
This analysis highlights the role of the Kreiss constant in capturing transient growth, particularly under non-normal dynamic conditions. While the spectral radius confirms asymptotic stability, the Kreiss constant reveals critical differences in transient behavior, providing insights into the system’s robustness and susceptibility to perturbations.
4.3 Floquet Analysis for Periodic Stability
[…]
4.4 Ablation Study and Benchmark
@@ 5 Conclusion
This paper introduces SALSA-RL, an RL framework for stability analysis in latent action space, leveraging action dynamics and deep learning. The utility of SALSA-RL is demonstrated across various environments, showcasing how control policies achieve stability, avoiding unstable regions, and maintaining stable states. Analysis of unstable regions reveals how key state-action combinations influence system stability, offering insights for improved control actions and reward function designs. Additional numerical analyses of periodic stability and transient growth provide a deeper understanding of underlying control outcomes, further interpreting the relationship between system states and stability.
Moreover, SALSA-RL’s latent linear dynamics yield coherent action regions in phase space, reflecting consistent actions in state space (see Appendix G). Finally, SALSA-RL performs competitively with popular DRL methods, while enhancing interpretability. It is therefore a powerful, yet simple tool for understanding the behavior of pretrained agents from various RL algorithms, offering insights into policy behavior, state transitions, and system stability.

### v3

@@ 3.5 Transient Growth Analysis
In the previous section, we used local stability analysis of latent (i.e., approximate) action dynamics as a proxy for assessing the local stability of a trained agent. However, even when the local stability condition $\rho(\mathbf{A}_{t})<1$ is met—non-normality (i.e., $\mathbf{A}_{t}\mathbf{A}_{t}^{*}\neq\mathbf{A}_{t}^{*}\mathbf{A}_{t}$) can still induce transient growth. This transient amplification, driven by the properties of $\mathbf{A}_{t}$, can destabilize intermediate states, amplify noise, or disrupt training. Due to the local nature of our stability analysis, the transient growth study becomes important for ascertaining whether a region in state-action phase space may ‘escape’ into a region of unstable action evolution.
For the latent dynamics described in Equation 4, the spectral radius $\rho(\mathbf{A}_{t})$ is first computed to confirm the system stability. Second, for cases where the system is locally stable ($\rho<1$) but $\mathbf{A}_{t}$ is non-normal, transient growth is analyzed using the Kreiss constant $\eta(\mathbf{A}_{t})$, which measures the potential short-term amplification of the system. The Kreiss constant is defined as:
@@
where $z$ is sampled from the complex plane outside the unit circle (i.e., $|z|>1$). A high $\eta(\mathbf{A}_{t})$ suggests a higher likelihood of transient growth. In such a case, one may observe an agent ultimately performing in an unstable manner even though the starting point of a trajectory has a spectral radius that is less than 1. Algorithm details are provided in Appendix B. Overall, the combination of local spectral radius computation and Kreiss constant analysis provides a qualitative understanding of the locally transient behavior of the system, in the absence of global stability analysis. This ensures an indirect assessment of the robustness of the learned control through our proposed framework.
3.6 Floquet Analysis for Periodic Stability
Floquet analysis [30] is a method specifically designed to evaluate the stability of periodic systems and their behavior over time. It quantifies the evolution of small perturbations in the latent space, capturing their growth, decay, or oscillatory behavior through Floquet exponents. This complements the spectral radius and Kreiss constant analyses by focusing on periodic stability, where transient growth alone may not capture recurring instabilities or oscillations. This analysis is applicable when the linear term in the latent dynamical system exhibits periodic trends, as seen in tasks requiring repeated oscillatory actions. An example is a pendulum failing to swing upright and instead oscillating at the bottom, where periodic stability is essential for sustained motion.
@@ 3.6 Floquet Analysis for Periodic Stability
In SALSA-RL framework, $\mathbf{A}_{t}$ is directly derived from the state $\mathbf{s}_{t}$, enabling the period to be identified by analyzing state information. Floquet exponents are computed at the end of each episode to diagnose latent dynamics, revealing instability, oscillations, or unstable growth. This analysis identifies regions of instability, oscillatory behavior, or unstable growth. An intuitive application for this is for periodic systems, such as for controlling the pendulum or cartpole benchmarks in RL. A step-by-step procedure is provided in Appendix B.
@@ 4 Experiments
A range of classical RL environments are selected from OpenAI Gym [5], focusing on continuous control tasks such as Cartpole-v1 (with continuous action space), Pendulum-v1, LunarLanderContinuous-v2, and BipedalWalker-v3. Additionally, a modified LunarLander is analyzed with a hovering objective to enhance stability and interpretability (details in Appendix D).
4.1 Eigenvalue-based Local Stability Analysis
@@
Pendulum. This task involves controlling a pendulum using a single torque input at the pivot. The objective is not just to swing the pendulum upright, but to position it in the upright position and maintain balance without further oscillation or swinging. This requires precise and continuous control to counteract gravity. A latent size of $d_{h}=3$ is used for model training and eigenvalue-based local stability analysis. Figure 2 shows the system’s behavior over time (as a trajectory) and the eigenvalue-based evaluation (underlying contour).
Figure 2: Local stability analysis of Pendulum control. The white line marks where the spectral radius equals 1, and the red line shows the observed state trajectory. Contours illustrate local variations in the norm behavior of latent dynamics, as captured by $\rho(\mathbf{A}_{t})$. The trajectory initially enters regions of local norm growth (i.e., $\rho>1$), but gradually shifts into regions of local contraction ($\rho<1$), effectively maintaining the system within a range that avoids high-growth zones (bounded by the white line).
@@
In the first column of the analysis, the trajectory (red line) evolves from locally unstable regions to stable ones. Initially, as the pendulum swings up from the bottom position ($\cos\theta,\sin\theta<0$), it exhibits instability, reflected by high spectral radius values ($\rho>1$) in the action space. Over time, the pendulum stabilizes in the upright position ($\cos\theta\to 1$, $\sin\theta\to 0$), with the spectral radius dropping below 1 ($\rho<1$), indicating entry into a stable regime and steady-state behavior. In the second column, the imaginary eigenvalues remain at 0, confirming local stability. While states initially approach regions with positive imaginary values, suggesting potential oscillations, they eventually shift away, guiding the system to a steady state.
In stabilized regimes, we observe that the latent state continues to evolve, i.e., $\mathbf{z}_{t+1}\neq\mathbf{z}_{t}$, despite the system being locally stable. This is due in part to the decode–re-encode cycle applied at each step: the predicted latent $\mathbf{z}_{t+1}$ is first decoded into an action $a_{t+1}=\mathrm{Decoder}(\mathbf{z}_{t+1})$, which is then re-encoded as $\mathbf{z}_{t+1}^{\prime}=\mathrm{Encoder}(a_{t+1})$ for the next timestep. The updated latent state then evolves via Equation 4 as $\mathbf{z}_{t+2}=(I+\mathbf{A}_{t})\mathbf{z}_{t+1}^{\prime}\approx\mathbf{z}_{t+1}$, causing stabilized decoded actions $a_{t+2}\approx a_{t+1}$.
This phenomenon arises from the non-unique nature of the encoder–decoder mapping, along with the fact that decoding and re-encoding are performed at every step (see Section 3.2 for details on the decoder module). As a result, while the latent representation may vary, the decoded actions remain effectively invariant across time steps. This practical invariance in the action space underscores the reliability of our latent local stability diagnostics: even as latent states drift, the resulting control behavior remains consistent, highlighting the model’s ability to stabilize behavior through structured latent dynamics.
@@
Additionally, Appendix E further examines regions of local instability, recovery, and control failures, reinforcing the framework’s stability analysis. Overall, this analysis shows the framework policy not only stabilizes the pendulum but does so in a manner that aligns with the primary objective of keeping it upright. By maintaining eigenvalues within the locally stable region and avoiding oscillatory dynamics, the system ensures stability while providing interpretable insights into how the policy achieves the desired control objectives.
CartPole. The CartPole environment involves a single discrete action that moves the cart left or right to balance the pole. With a latent dimension of $d_{h}=3$, the trained model effectively learns to stabilize the pole in the upright position, as shown in Figure 3.
Figure 3: Local stability analysis of CartPole control. Each row shows a representative state. Due to the nature of the task, the system frequently operates in regions with $\rho(\mathbf{A}_{t})>1$, as the cart must oscillate to keep the pole upright. These excursions near or beyond the contraction boundary reflect the need for continuous corrective actions rather than convergence to a fixed point. Thus, regions with $\rho(\mathbf{A}_{t})>1$ are not necessarily undesirable, but rather reflect the non-equilibrium behavior required by the task.
However, the spectral radius tends to fluctuate near the white line $\rho=1$, even over extended periods. This behavior can be attributed to the fact that the cart must constantly move left or right to counteract the pole’s dynamics. These continual adjustments introduce small, persistent deviations in the latent dynamics, keeping the system near the edge of stability despite overall successful control. Meanwhile, the imaginary eigenvalue contours reveal a similar pattern of oscillatory behavior. The values consistently fluctuate between zero and nonzero, indicating persistent but bounded oscillations in the latent dynamics throughout the episode.
@@
LunarLander (Hovering Objective). For Lunarlander experiment, we utilize a modified environment with a hovering objective instead of an objective to land safely as typically utilized. Since the state includes a 2D space, eigenvalue-based local stability is evaluated across the entire 2D domain, with each frame in Figure 4 representing a single timestep.
@@
Figure 4: Local stability analysis of LunarLander hovering control. The white line indicates where spectral radius $\rho$ equals 1, separating regions of local stability and instability. The lander is initially deployed in various regions beyond the default upper-middle position (case 1). Each frame shows eigenvalue distributions in 2D space, highlighting stability regions, with the trajectory about 1,000 timesteps. Notably, case 2 contours provide early warnings of instability before the lander crashes. Animations are provided in the Supplementary.
@@
Among Cases 1, 3, and 4, the spectral radius consistently transitions from high-value regions to lower-value regions, reflecting a transition from locally unstable dynamics to more stable behavior over time. This gradual shift is visually indicated by the trajectory moving from red/yellow zones to cooler regions in the contour plots, especially in the later timesteps. The imaginary contour highlights regions of oscillatory behavior, and the lander’s actual trajectory exhibits a transition from high to low oscillation over time. In case 2, however, the lander crashes at the final frame. The spectral radius contour in this case reveals large regions of local instability, aligning with the lander’s inability to maintain its hovering status. Notably, the contour shows that the current state-action combination enters a highly unstable region, effectively signaling the risk of a crash before it occurs.
In SALSA-RL, the latent action is not clipped; the clamp is applied only after decoding. Detecting latent-space instability, therefore, pinpoints where the policy relies on saturation rather than manual clipping, as unbounded latent growth forces the decoder to flip between clipped extremes. In this modified LunarLander task, spikes in latent-action instability emerged several steps before any visible state deviation and reliably predicted impending crashes, even though the thrust values themselves remained within their clipped limits.
Overall, this strongly underscores the value of our local stability analysis, as it provides early warnings of imminent failures. Because action stability ultimately drives state evolution, latent action-space analysis yields policy-level interpretability, showing how the agent reacts over time rather than merely its eventual state. Empirically, in the modified LunarLander task, spikes in latent-action instability consistently preceded crashes even while the visible state trajectory still looked nominal. Overall, SALSA-RL is not a replacement for state-based analysis but a complementary, policy-aware diagnostic tool.
Besides the above 3 experiments, Appendix F explores the BipedalWalker benchmark with additional results and discussions, reinforcing the framework and analysis. Furthermore, Appendix I extends SALSA-RL to the high-dimensional Humanoid environment, where it achieves performance comparable to PPO while additionally enabling post-hoc stability analysis. These results demonstrate the scalability and robustness of the framework, highlighting how the latent-space structure supports local stability analysis and offers increasing benefits in higher-dimensional or more complex control tasks.
4.2 Transient Growth Analysis
Extending the local stability analysis, the Kreiss constant $\eta$ is evaluated to assess transient growth in the pendulum under a standard setup (Case 1) and a modified environment (Case 2), where gravity increases from 10 to 15 and mass from 1.0 to 1.1. The third row of Figure 5 illustrates $\eta$ over time for both cases.
Figure 5: Transient Growth and Floquet analysis of Pendulum control. The pendulum achieves stabilization in case 1 (first column) while displaying periodic behavior with oscillations in case 2 (second column). The Kreiss constant (third row) shows differing transient growth behaviors over time. The resultant Floquet exponents are $\mu_{i}=0$ and $\mu_{i}=\bigl{\{}0.96,\ 0.02\pm 0.04i\bigr{\}}$ for cases 1 and 2, respectively.
In Case 1, the pendulum successfully stabilizes in the upright position. This is reflected in the actions and states, which converge to near-zero values without oscillations (first and second rows). The Kreiss constant remains small throughout, indicating minimal transient growth and robust stability.
In Case 2, the increased gravity and mass disrupt the stabilization process, causing the pendulum to exhibit periodic motion instead of reaching the upright position. This behavior is evident in the actions and states, which display sustained oscillations over time. The Kreiss constant shows periodic spikes, reflecting transient growth that aligns with the observed oscillatory response. These spikes indicate significant susceptibility to transient amplification, even though the system remains bounded over time.
This highlights the role of the Kreiss constant in capturing transient growth, particularly under non-normal dynamics. While spectral radius confirms asymptotic stability, the Kreiss constant reveals critical differences in transient behavior, providing insights into the system’s robustness and susceptibility to perturbations.
4.3 Floquet Analysis for Periodic Stability
In Case 1 (left column of Figure 5), the Floquet analysis reveals all exponents $\mu_{i}=0$ across the three-dimensional latent space, indicating marginal stability. This means that small perturbations neither grow nor decay over the period, and the system maintains its oscillatory trajectory without exponential divergence or damping. Zero imaginary components further confirms that the observed oscillations are structural or externally driven, rather than intrinsic to the latent dynamics.
In Case 2 (right column of Figure 5, modified environment), the $\mu_{i}=\bigl{\{}0.96,\ 0.02\pm 0.04i\bigr{\}}$ reveal instability, with positive real parts driving perturbation growth and imaginary components introducing oscillatory modes. These dynamics correspond to the pendulum’s observed periodic motion, characterized by high-frequency control oscillations, significant angular velocity peaks, and sustained oscillatory state trajectories.
4.4 Ablation Study and Benchmark
@@ 5 Conclusion
This paper introduces SALSA-RL, an RL framework for local stability analysis in latent action space, leveraging action dynamics and deep learning. The utility of SALSA-RL is demonstrated across various environments, showcasing how control policies achieve stability, avoiding unstable regions, and maintaining stable states. Analysis of unstable regions reveals how key state-action combinations influence system stability, offering insights for improved control actions and reward function designs. Additional numerical analyses of periodic stability and transient growth provide a deeper understanding of underlying control outcomes, further interpreting the relationship between system states and stability.
Moreover, SALSA-RL’s latent linear dynamics yield coherent action regions in phase space, reflecting consistent actions in state space (see Appendix G). Importantly, the stability analysis module functions as a post-hoc diagnostic: it is applied after training without interfering with optimization, making the approach algorithm-agnostic in principle and broadly applicable across DRL methods. SALSA-RL may also complement safe and stability-focused RL approaches [15, 6, 17, 61, 22], by offering post-hoc analysis tools to interpret policy behavior without modifying the training process.
Finally, the goal of SALSA-RL is not to outperform on saturated benchmarks but to introduce an interpretable framework for empirical stability analysis of RL policies—capabilities typically lacking in existing RL algorithms. It is therefore a powerful yet simple tool for understanding the behavior of pretrained agents from various RL algorithms, offering insights into policy behavior, state transitions, and system stability. Nonetheless, SALSA-RL currently focuses on post-hoc stability analysis and does not enforce safety or correct instability during training. Extending the framework to actively guide policy updates—potentially in conjunction with safe RL methods—remains an important direction for future work.
6 Acknowledgements
This research used resources of the Argonne Leadership Computing Facility (ALCF), a U.S. Department of Energy (DOE) Office of Science user facility at Argonne National Laboratory, and is based on research supported by the U.S. DOE Office of Science-Advanced Scientific Computing Research Program, under Contract No. DE-AC02-06CH11357. RM and XL acknowledge funding support from the DOE Office of Science, Fusion Energy Sciences program (DOE-FOA-2905, PM-Dr. Michael Halfmoon), and computational resources from the Penn State Institute for Computational and Data Sciences.